Calculate the pH of a solution that was prepared by mixing 25.00 mL of a 0.1630 M HClO4 solution with 20.00 mL of a 0.1350 Ba(OH)2 solution


1
Expert's answer
2021-05-26T07:04:25-0400

pH is a parameter which is used to determined the acidity or basicity of an aqueous solution. 

pH is determined by the concentration of either H+or OH- ion concentration.

pH = -log[H+]

pOH = -log[OH-]


HClO4 Strong Acid

V1=25mL

M1=0.1630M

n1=1(total change of a cation or anion)

Ba(OH)2 strong base

V2=20mL

M2=0.1350M

n2=2

Ba2++2(OH-)


N.o. of H+ ion In HClO4-


n1M1V1=1×25×0.1630=4.075


N. o. Of OH- ion from Ba(OH)2

n2M2V2=2×20×0.1350=5.4


Hence n2M2V2>n1M1V1


[OH-]= "\frac{5.4-4.075}{20+25}"


"=0.02944"


pOH=-log[OH-]

=-log(0.02944)

pOH =1.53

But pH=14-pOH

=14-1.53

=12.47

pH=12.47